# Highly accurate simulations of asymmetric black-hole scattering and cross validation of effective-one-body models

### v1

@@ IV.3 Unequal-mass non-spinning black holes
The left panel of Fig. 10 presents the scattering angle as a function of the symmetric mass ratio for simulations with $\gamma=1.02$ and $\ell=4.8$. The upper plot shows that while a decrease in the symmetric mass ratio does slightly increase the scattering angle, the change across the range of mass ratios is small. The lower panel shows that all of the EOB models capture this behaviour well and are all within less than a percent of the NR results across all the mass ratios. One interesting thing to note is that the SEOBNRv5and SEOB-4PM models tend towards each other as the mass ratio increases ($\nu$ decreases). This is due to the fact that both models are designed to recover the test mass limit.
The right panel of Fig. 10 shows the scattering angle for simulations with $\gamma=1.0916$ and $b=10.5M$. Interestingly the behaviour for these parameters is very different from the one just discussed. The upper plot shows that, as before, a decreasing symmetric mass ratio increases the scattering angle, but the change is very significant, including a capture at $q=8$ ($\nu\approx 0.1$). Both evolution models SEOBNRv5 and TEOBResumS-Dalí fail to accurately predict the capture with the former predicting a capture at $q\approx 10$ and the latter predicting a capture at $q\approx 5$. The lower panel shows that most of the EOB models are primarily within a few percent of the NR results at comparable masses but diverge as the symmetric mass ratio decreases. The noteable exception is the TEOBResumS-Dalí model which shows worse agreement across the whole range of mass ratios. These increased differences across all models are likely due to the fact that we are approaching the separatrix and thus in the strong-field regime where we expect the accuracy of all the models to decrease.
Figure 11: Scattering angle results for unequal mass non-spinning systems with constant ($\gamma\!=\!1.02$, $\ell\!=\!4.8$) (left) and ($\gamma\!=\!1.0916$, $b\!=\!10.5M$) (right). $\theta_{1}$ ($\theta_{2}$) represents the scattering angle of the larger (smaller) BH. The difference in the scattering angles is shown in the lower panel. The vertical dotted line on the right plot shows the first confirmed capture of the SpEC simulations.
@@
Figure 11 also shows the estimates of each BH angle from the SEOB-PM model. The lower energy dataset (left) shows that both models follow the same general shape of the NR data, with zero difference in the angles at equal mass, followed by the difference rising quickly to a peak and then tending towards equal angles in the test body limit. However, when comparing to the NR data, the 4PM model underestimates the difference in the angles whereas the addition of the 5PM(1SF) terms causes the model to overestimate. The higher energy dataset (right) shows very different behaviour. Both models have the zero difference in the equal mass limit but then quickly diverge as the mass ratio increases. We hypothesise that this is due to the fact that we are approaching the separatrix, beyond which the notion of two scattering angles does not make sense.
@@ V Conclusions
The key finding of this study is the first measurement of disparate scattering angles from NR simulations due to the asymmetric emission of GWs. We have shown that this effect, though small ($\lesssim 1\%$ of the total angle), is extractable from SpEC simulations for both anti-parallel spin systems and unequal mass systems. We also present the first closed-form models constructed to calculate the scattering angle from a single BH, rather than the relative angle of both BHs. Comparing to the NR results shows that these SEOB-PM models struggle to accurately predict the difference in the scattering angles due to it being a high-order perturbative effect.
Having done the groundwork for NR scatter simulations from SpEC, it would be an exciting future perspective to extend the parameter space. Here, a particularly intriguing direction would be scattering orbits in the high-energy limit. Importantly, in this limit, the PM regime is no longer valid and new theoretical approximations would be required to inform the EOB models. Namely, as already noticed by Kovacs and Thorne Kovacs and Thorne (1978), perturbative scattering results are only valid provided that the overall deflection angle is small compared with the inverse boost parameter $\gamma^{-1}$ which in the high-energy limit implies:
@@ Appendix A Effect of junk radiation
In order to use these formulae to estimate the effect of the junk radiation, we perform a series of trial runs with different initial separations and analyse the differences in the post-junk masses and trajectories. The results of these simulations are shown in Fig. 12 for a set of simulations with $(\Gamma,\ell,q)=(1.02264,4.5367,1)$ that start at initial separations $D_{0}=(175,250,350)$. The top panel of Fig. 12 shows the relative velocity of the two BHs as a function of separation on the ingoing leg. Each curve shows an initial pulse at the start of the simulation as well as a secondary pulse when the junk radiation emitted from one BH interacts with the other. The middle and bottom panels show the difference in the energy and angular momentum of the simulations from Eqs. (33) and (34). The curves show that the differences between all of the configurations decay after the initial pulses to reach discrepancies of $\lesssim 10^{-5}$.
@@
Appendix B Scattering angle values
@@
$-0.26$ $-0.26$ $405.019$ $510.560$
$-0.25$ $-0.25$ $385.441^{+0.020}_{-0.047}$ $380.369$ $437.515$
$-0.20$ $-0.20$ $300.868^{+0.009}_{-0.005}$ $310.446$ $323.630$
$-0.10$ $-0.10$ $238.782^{+0.007}_{-0.005}$ $247.779$ $251.214$
$-0.00$ $0.00$ $207.041^{+0.005}_{-0.004}$ $214.149$ $216.177$ $226.521$ $319.800$
$0.10$ $0.10$ $186.111^{+0.005}_{-0.005}$ $191.701$ $193.518$ $172.404$ $238.814$
$0.20$ $0.20$ $170.712^{+0.007}_{-0.008}$ $175.156$ $177.033$ $151.030$ $203.180$
$0.30$ $0.30$ $158.656^{+0.009}_{-0.011}$ $162.224$ $164.215$ $137.689$ $181.149$
$0.40$ $0.40$ $148.875^{+0.010}_{-0.013}$ $151.717$ $153.814$ $127.987$ $165.616$
$0.50$ $0.50$ $140.713^{+0.012}_{-0.016}$ $142.941$ $145.113$ $120.399$ $153.841$
$0.60$ $0.60$ $133.738^{+0.014}_{-0.019}$ $135.508$ $137.725$ $113.860$ $144.609$
Table 3: Scattering angles (in degrees) for equal mass BHs with spin aligned with the orbital angular momentum. All of these simulations have constant $\Gamma=1.055$ and $\ell=5.18$. A dash (—) indicates a capture.
@@
Table 6: Scattering angles (in degrees) for unequal mass, non-spinning BHs with $\gamma\!=\!1.0916$ and $b\!=\!10.5M$. SpEC values are given for both BHs, with $\theta_{1}$ corresponding to the larger mass. EOB values correspond to the relative scattering angle. A dash (—) indicates a capture.

### v2

@@ IV.3 Unequal-mass non-spinning black holes
The left panel of Fig. 10 presents the scattering angle as a function of the symmetric mass ratio for simulations with $\gamma=1.02$ and $\ell=4.8$. The upper plot shows that while a decrease in the symmetric mass ratio does slightly increase the scattering angle, the change across the range of mass ratios is small. The lower panel shows that all of the EOB models capture this behavior well and are all within less than a percent of the NR results across all the mass ratios with $w_{4{\rm PM}}$ showing the best agreement throughout. One interesting thing to note is that the SEOBNRv5and SEOB-4PM models tend towards each other as the mass ratio increases ($\nu$ decreases). This is due to the fact that both models are designed to recover the test mass limit.
The right panel of Fig. 10 shows the scattering angle for simulations with $\gamma=1.0916$ and $b=10.5M$. Interestingly the behavior for these parameters is very different from the one just discussed. The upper plot shows that, as before, a decreasing symmetric mass ratio increases the scattering angle, but the change is very significant, including a capture at $q=8$ ($\nu\approx 0.1$). Both evolution models SEOBNRv5 and TEOBResumS-Dalí fail to accurately predict the capture with the former predicting a capture at $q\approx 10$ and the latter predicting a capture at $q\approx 5$. The lower panel shows that most of the EOB models are primarily within a few percent of the NR results at comparable masses but diverge as the symmetric mass ratio decreases. The noteable exception is the TEOBResumS-Dalí model which shows worse agreement across the whole range of mass ratios. These increased differences across all models are likely due to the fact that we are approaching the separatrix and thus in the strong-field regime where we expect the accuracy of all the models to decrease.
Figure 11: Scattering angle results for unequal mass non-spinning systems with constant ($\gamma\!=\!1.02$, $\ell\!=\!4.8$) (left) and ($\gamma\!=\!1.0916$, b=10.5Mb\!=\!10.5M) (right). $\theta_{1}$ ($\theta_{2}$) represents the scattering angle of the larger (smaller) BH. The difference in the scattering angles is shown in the lower panel. The vertical dotted line on the right plot shows the first confirmed capture of the SpEC simulations. In these comparisons, we use the asymmetrical closed-form SEOB-PM models based on the individual scattering angle $\theta_{i}$.
@@
Figure 11 also shows the estimates of each BH angle from the asymmetrical SEOB-PM model based on the individual scattering angles $\theta_{i}$. The lower energy dataset (left) shows that both models follow the same general shape of the NR data, with zero difference in the angles at equal mass, followed by the difference rising quickly to a peak and then tending towards equal angles in the test body limit. However, when comparing to the NR data, the 4PM model underestimates the difference in the angles whereas the addition of the 5PM(1SF) terms causes the model to overestimate. The higher energy dataset (right) shows very different behavior. Both models have the zero difference in the equal mass limit but then quickly diverge as the mass ratio increases. We hypothesize that this is due to the fact that we are approaching the separatrix, beyond which the notion of two scattering angles does not make sense.
@@ V Conclusions
The key result of this study is the first measurement of disparate scattering angles from NR simulations due to the asymmetric emission of GWs. We have shown that this effect, though small ($\lesssim 1\%$ of the total angle), is extractable from SpEC simulations for both anti-parallel spin systems and unequal mass systems. We also present the first closed-form models constructed to calculate the scattering angle from a single BH, rather than the relative angle of both BHs. Comparing to the NR results shows that these SEOB-PM models struggle to accurately predict the difference in the scattering angles due to it being a high-order perturbative effect.
Having done the groundwork for NR scatter simulations from SpEC, it would be an exciting future perspective to extend the parameter space. Here, a particularly intriguing direction would be scattering orbits in the high-energy limit. Importantly, in this limit, the PM regime is no longer valid and new theoretical approximations would be required to inform the EOB models. Namely, as pointed out by D’Eath D’Eath (1978) and then Kovacs and Thorne Kovacs and Thorne (1978), perturbative scattering results are only valid provided that the overall deflection angle is small compared with the inverse boost parameter $\gamma^{-1}$ which in the high-energy limit implies:
@@ Appendix A Effect of junk radiation
In order to use these formulae to estimate the effect of the junk radiation, we perform a series of trial runs with different initial separations and analyze the differences in the post-junk masses and trajectories. The results of these simulations are shown in Fig. 12 for a set of simulations with $(\Gamma,\ell,q)=(1.02264,4.5367,1)$ that start at initial separations $D_{0}=(175,250,350)$. The top panel of Fig. 12 shows the relative velocity of the two BHs as a function of separation on the ingoing leg. Each curve shows an initial pulse at the start of the simulation as well as a secondary pulse when the junk radiation emitted from one BH interacts with the other. The middle and bottom panels show the difference in the energy and angular momentum of the simulations from Eqs. (33) and (34). The curves show that the differences between all of the configurations decay after the initial pulses to reach discrepancies of $\lesssim 10^{-5}$.
@@
Appendix B Details of the SEOB-5PM(1SF) model
As discussed in the main text, the PM content of the closed-form models may be characterized by their accuracy after PM expanding both sides of Eq. (22) which we reprint here for clarity:
$\theta+\pi=-2\int_{r_{\rm min}}^{\infty}\!{\rm d}r\frac{\partial}{\partial L}p_{r}\,.$ (35)
The PM expansion of the right-hand-side is non-trivial but discussed e.g. in Refs. Damour and Schaefer (1988); Damour (2020); Buonanno et al. (2024b). For the nnPM closed-form models with $n\leq 4$ the PM expansion of the the right-hand-side of Eq. (35) is simply required to match the analytic nnPM angle of the left-hand-side.
As stated in the main text, the 5PM angle has a 2SF component which is not known. Thus, in the PM matching of the SEOB-5PM(1SF) model, the $G^{5}$ part of the relative scattering angle is accurate only to 1SF order. More precisely, using the notation of Ref. Driesse et al. (2025), the analytic angle $\theta^{\prime}_{\rm rel}$ used for matching is,
$\theta^{\prime}_{\rm rel}\!$ $=\!\Gamma\sum_{n=1}^{4}\bigg(\!\frac{GM}{b}\!\bigg)^{n}\sum_{m=0}^{\lfloor\frac{n-1}{2}\rfloor}\nu^{m}\theta^{(n,m)}+\frac{1}{\Gamma}\bigg(\!\frac{GM}{b}\!\bigg)^{4}\nu^{2}\theta^{(4,2)}$
$+\Gamma\bigg(\!\frac{GM}{b}\!\bigg)^{5}\Big[\theta^{(5,0)}+\nu\theta^{(5,1)}+\nu^{2}\frac{2+53v^{2}}{5v^{6}}\Big]\ ,$ (36)
where $v=\sqrt{\gamma^{2}-1}/\gamma$ is the relative velocity. The first line of Eq. (36) is the full 4PM relative angle while the next line gives the 1SF-accurate 5PM angle with additional 3PN-2SF terms. We neglect spin because the asymmetric SEOB-5PM(1SF) model is used only for non-spinning comparisons, though spin effects would not add any further subtleties. As discussed in the main text, the 3PN-2SF terms are included for self-consistency of the model. Namely, those terms are determined from the 4PM input following the usual correspondence that nnPM determines (n1)(n-1)PN. If one would not include the 3PN-2SF terms in the matching angle $\theta^{\prime}_{\rm rel}$, the 5PM(1SF) deformation would pick up unphysical poles in the $v\to 0$ limit.
Finally, for the asymmetric SEOB models, we compute the individual scattering angle using the following relation,
$\theta_{1}=\theta_{\rm rel}-\frac{E_{1}}{M\mu\sqrt{\gamma^{2}-1}}(\hat{\bm{b}}-\theta_{\rm rel}\hat{\bm{p}})\cdot\bm{P}_{\rm recoil}+\mathcal{O}(G^{6})\ ,$ (37)
which is similar to Eq. (27) and uses the same notation. An analogous equation for $\theta_{2}$ is obtained by symmetry. For the SEOB-5PM(1SF) one requires the 5PM piece of the $\hat{\bm{b}}-$direction of the recoil which is 1SF-exact (see Ref. Driesse et al. (2025)). The matching angle $\theta^{\prime}_{1}$ for the asymmetric SEOB-5PM(1SF) model is then computed using Eq. (37) with $\theta_{\rm rel}$ given by $\theta_{\rm rel}^{\prime}$ defined above and the recoil from Ref. Driesse et al. (2025) (throwing away any terms of order $G^{6}$ or beyond).
Finally, we note that the $G^{5}$ contributions were included in a PN-expanded manner. Namely, they were expanded to $\mathcal{O}(v^{25})$, corresponding to 15PN order. In the future it would be interesting to implement direct numerical evaluation of the multiple polylogarithms of maximal weight three required for these $G^{5}$ contributions.
Appendix C Scattering angle values
@@
0.26-0.26 0.26-0.26 $405.019$ $510.560$
0.25-0.25 0.25-0.25 $385.441^{+0.020}_{-0.047}$ $380.369$ $437.515$
0.20-0.20 0.20-0.20 $300.868^{+0.009}_{-0.005}$ $310.446$ $323.630$
0.10-0.10 0.10-0.10 $238.782^{+0.007}_{-0.005}$ $247.779$ $251.214$
$\phantom{-}0.00$ $\phantom{-}0.00$ $207.041^{+0.005}_{-0.004}$ $214.149$ $216.177$ $226.521$ $319.800$
$\phantom{-}0.10$ $\phantom{-}0.10$ $186.111^{+0.005}_{-0.005}$ $191.701$ $193.518$ $172.404$ $238.814$
$\phantom{-}0.20$ $\phantom{-}0.20$ $170.712^{+0.007}_{-0.008}$ $175.156$ $177.033$ $151.030$ $203.180$
$\phantom{-}0.30$ $\phantom{-}0.30$ $158.656^{+0.009}_{-0.011}$ $162.224$ $164.215$ $137.689$ $181.149$
$\phantom{-}0.40$ $\phantom{-}0.40$ $148.875^{+0.010}_{-0.013}$ $151.717$ $153.814$ $127.987$ $165.616$
$\phantom{-}0.50$ $\phantom{-}0.50$ $140.713^{+0.012}_{-0.016}$ $142.941$ $145.113$ $120.399$ $153.841$
$\phantom{-}0.60$ $\phantom{-}0.60$ $133.738^{+0.014}_{-0.019}$ $135.508$ $137.725$ $113.860$ $144.609$
Table 3: Scattering angles (in degrees) for equal mass BHs with spin aligned or anti-aligned with the orbital angular momentum. All of these simulations have constant $\Gamma=1.055$ and $\ell=5.18$. A dash (—) indicates a capture.
@@
Table 6: Scattering angles (in degrees) for unequal mass, non-spinning BHs with $\gamma\!=\!1.0916$ and b=10.5Mb\!=\!10.5M. SpEC values are given for both BHs, with $\theta_{1}$ corresponding to the larger mass. EOB values correspond to the relative scattering angle. A dash (—) indicates a capture.